# Cubic fourfolds with symplectic automorphisms

### v1

@@ 1. Introduction
More recently, Yang, Yu and Zhu studied full automorphism groups for cubic fivefolds and fourfolds in [YYZ23]. According to them, there are 15 maximal groups, and a finite group $G$ can acts faithfully on a smooth cubic fourfold if and only if $G$ is a subgroup of these 15 groups. However, explicit equatoins of cubic fourfolds with automorphisms are still unknown in many cases.
In this paper, we study symplectic actions of 34 groups of Laza and Zheng, and detertmine invariant smooth cubic fourfolds. Since any automorphism of a smooth cubic fourfold is given by a projective transformation, our problem is reduced to classification of faithfull projective representations of 34 groups which satisfy symplectic conditions. It is known that complex projective representations of a finite group $G$ lift to linear representations of a Schur cover $G^{*}$. We determine possible projective actions of 34 groups $G$ by considering character tables of $G^{*}$. As the result, we have
@@
[…]
For other cases, the moduli spaces are irreducible (see Table 1 in sectoin 2 ).
For maximal cases, defining eqautions of corresponding cubic fourfolds are known, and they are defined over $\mathbb{Q}$ in most cases. In the last section, we show
@@
[…]
@@
  • $p^{1+2n}_{\pm}$ denote extraspecial groups of order $p^{1+2n}$.
@@ 2.1. Schur covering groups
and the symmetric group $\mathrm{S}_{7}$ acts on it as permutations of coordinates (However, the actoin of $(12)\in\mathrm{S}_{7}$ is not symplectic). Another $\mathrm{A}_{7}$-invariant cubic fourfold was identified by Yang, Yu and Zhu in [YYZ23] considering the “triple cover” of $\mathrm{A}_{7}$. Their model is defined over $\mathbb{Q}(\sqrt{15})$, and we give an equation over $\mathbb{Q}$ here. Let $G$ be a subgroup of $\mathrm{GL}_{6}(\mathbb{C})$ generated by two matrices
@@
[…]
such that $M\cong\mathrm{H}^{2}(G,\mathbb{C}^{\times})$ and $\iota(M)\subset[G^{*},G^{*}]$. In general, $G^{*}$ is not uniquely determined. The Klein’s four group $\mathrm{C}_{2}\times\mathrm{C}_{2}$ has two Schur covering groups $\mathrm{D}_{8}$ and $\mathrm{Q}_{8}$. However, it is known that all projective representations of $G$ can be lifted to linear representations of one Schur covering group $G^{*}$. An abelian group $M=M(G)$ is called the Schur multiplier of $G$. For example, we have
@@
(see [Ka85]). If $M(G)=1$, we have $G^{*}=G$ and any projective representation of $G$ is lifted to a linear representation of $G$ itself.
@@
It is known that $G$ has only one Schur covering group if the orders of $G/[G,G]$ and $M(G)$ are coprime ([Ka85], Theoerm 4.4).
For non-maximal cases of Laza and Zheng’s 34 groups, we have the following table:
Table 1. Schur covers and multipliers
$r(G)$ $G$ GAP-ID $M$ ID* Eq irr
8 2 2,1 1 2,1 (3.1) 1
12 $2^{2}$ 4,2 2 8,3 (3.2) 1
12 3 3,1 1 3,1 (3.3), (3.4) 2
14 4 4,1 1 4,1 (3.5) 1
14 $\mathrm{S}_{3}$ 6,1 1 6,1 (3.6) 1
15 $\mathrm{D}_{8}$ 8,3 2 16,7 (3.7) 1
16 $\mathrm{A}_{3,3}$ 18,4 3 54,8 (4.6), (4.11) 2
16 $\mathrm{D}_{12}$ 12,4 2 24,8 (4.2), (4.3) 2
16 $\mathrm{A}_{4}$ 12,3 2 24,3 (4.20), (4.22) 2
16 $\mathrm{D}_{10}$ 10, 1 1 10, 1 (4.1) 1
17 $\mathrm{S}_{4}$ 24,12 2 48,28 (5.2), (5.3) 1
17 $\mathrm{Q}_{8}$ 8,4 1 8,4 (5.1) 1
18 $23^{1+4}_{+}:$ 2 486, 249 $3^{5}$ (6.1) 1
18 $\mathrm{A}_{4,3}$ 72,43 6 432,258 (6.5), (6.7) 2
18 $\mathrm{A}_{5}$ 60, 5 2 120,5 (6.9), (6.10) 2
18 $43^{2}.$4 36,9 3 108,15 (6.11), (6.12) 2
18 $\mathrm{S}_{3,3}$ 36,10 2 72,22 (6.13) 1
18 $\mathrm{F}_{21}$ 21,1 1 21,1 (6.2) 1
18 $\mathrm{Hol}(5)$ 20, 3 1 20,3 (6.15) 1
18 $\mathrm{QD}_{16}$ 16,8 1 16,8 (6.3) 1
19 $2.23^{1+4}_{+}:$ 2.2 972, 776 $3^{4}$ (7.1) 1
19 $\mathrm{A}_{6}$ 360,118 6 - (2.6), (2.9) 2
19 $\mathrm{L}_{2}(7)$ 168,42 2 336,114 (7.2) 1
19 $\mathrm{S}_{5}$ 120,34 2 240, 89 (7.3), (7.6) 2
19 $\mathrm{M}_{9}$ 72,41 3 216,88 (7.9) 1
19 $\mathrm{N}_{72}$ 72,40 2 144,117 (7.10) 1
19 $\mathrm{T}_{48}$ 48,29 1 48,29 (7.12) 1
where $r(G)=\mathrm{rank}\ S_{G}$ (see [LZ22]), GAP-ID is “IdGroup($G$)” in GAP, $M$ is the Schur multiplier of $G$, and ID* is GAP-ID of a Schur covering group $G^{*}$ given by “SchurCover($G$)” in GAP. Equations of corresponding cubic 4-folds are given in Eq, and irr is the number of irreducible components of the moduli space $\mathcal{M}(G)$ of cubic fourfolds $X$ such that $G\subset\mathrm{Aut}^{s}(X)$.
@@ 2.2. Symplectic automorphisms of cubic 4-folds
$1,2,3,4,5,6,7,8,9,11,12,15$
If a cubic fourfold $X$ has a symplectic automorphism of oder $n=9,11,12,15$, then we have
@@
and $g$ is conjugate to $\mathrm{diag}(1,1,1,1,-1,-1)$ in $\mathrm{PGL}_{6}(\mathbb{C})$. Also note that, although we can not define the trace of $g\in\mathrm{PGL}_{6}(\mathbb{C})$, the absolute value $|\mathrm{Tr}\,\tilde{g}|$ for a lift $\tilde{g}\in\mathrm{GL}_{6}(\mathbb{C})$ does not depend on a lifting if $\tilde{g}$ is of finite order. These values are usefull information to classify symplectic actions from character tables.
@@ 2.3. Invariants
If $\chi$ is trivial, we say that $F(x)$ is a $\rho(G)$-invariant.  We denote the vectore space of $\rho(G)$-semi-invariants of degree $d$ with character $\chi$, by $V_{d}(\rho,\chi)$. If $\chi$ is trivial, it is denoted by $V_{d}(\rho)$.
@@
(3)  Let us assume that $d$ and $|\widehat{G}|$ are co-prime to each other. Then, for any $\chi\in\widehat{G}$, there exist $\theta\in\widehat{G}$ such that
@@ 2.4. Moduli spaces
$\mathrm{C}_{\rho}(G)$ =$=\{h\in\mathrm{GL}_{n}(\mathbb{C})\ |\ hgh^{-1}=g\ \ \text{for}\ \ {}^{\forall}g\in\rho(G)\},$
$\mathrm{N}_{\rho}(G)$ =$=\{h\in\mathrm{GL}_{n}(\mathbb{C})\ |\ h\rho(G)h^{-1}=\rho(G)\}.$
@@
$\mathrm{N}_{\rho,\chi}(G)$ =$=\{h\in\mathrm{N}_{\rho}(G)\ |\ \chi_{\rho}(hgh^{-1})=\chi_{\rho}(g)\ \ \text{for}\ \ {}^{\forall}g\in\rho(G)\}$
@@
Now let $G$ be one of Laza and Zheng’s 34 groups, $G^{*}$ be a Schur cover and $\rho:G^{*}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ be a lifting of a faithful projective actoin of $G$ on $\mathbb{P}^{5}$. If $\rho(G^{*})$ acts simplectically on a smooth cubic fourfold $F(x)=0$ with $F\in V_{3}(\rho,\,\chi)$, we have
@@
Note that we have $\dim\mathrm{N}_{\rho,\chi}(G^{*})\geq 1$ since ×$\mathbb{C}^{\times}\subset\mathrm{N}_{\rho,\chi}(G^{*})$. Therefore we have the following criterion to exclude unnecessary cases.
@@ 2.5. $\mathrm{A}_{6}$-invariant cubic fourfolds
$\begin{array}[]{c|rrrrrrrrrrrrrrrrr}\hline\cr\rm class&\rm 1&\rm 3A&\rm 3B&\rm 2A&\rm 6A&\rm 6B&\rm 3C&\rm 3D&\rm 12A&\rm 12B&\rm 4A&\cdots\\ \rm size&1&1&1&45&45&45&120&120&90&90&90&\cdots\\ \hline\cr\chi_{1}&1&1&1&1&1&1&1&1&1&1&1&\cdots\\ \chi_{2}&3&3\overline{\omega}&3\omega&-1&-\overline{\omega}&-\omega&0&0&\overline{\omega}&\omega&1&\\ \chi_{3}&3&3\overline{\omega}&3\omega&-1&-\overline{\omega}&-\omega&0&0&\overline{\omega}&\omega&1&\\ \chi_{4}&3&3\omega&3\overline{\omega}&-1&-\omega&-\overline{\omega}&0&0&\omega&\overline{\omega}&1&\\ \chi_{5}&3&3\omega&3\overline{\omega}&-1&-\omega&-\overline{\omega}&0&0&\omega&\overline{\omega}&1&\\ \chi_{6}&5&5&5&1&1&1&-1&2&-1&-1&-1&\\ \chi_{7}&5&5&5&1&1&1&2&-1&-1&-1&-1&\\ \chi_{8}&6&6\overline{\omega}&6\omega&2&2\overline{\omega}&2\omega&0&0&0&0&0&\\ \chi_{9}&6&6\omega&6\overline{\omega}&2&2\omega&2\overline{\omega}&0&0&0&0&0&\\ \vdots&\vdots&\\ \hline\cr\end{array}$
@@
Classes 1, 3A and 3B form the center $\mathrm{C}_{3}$, and $\chi_{i}\ (i=1,6,7)$ are irreducible characters with $\mathrm{C}_{3}\subset\mathrm{Ker}\,\chi_{i}$, that is, characters of $\mathrm{A}_{6}$ itself rather than the triple cover $3.\mathrm{A}_{6}$. In this case, we have two (non-trivial) characters $\chi_{1}+\chi_{6}$ and $\chi_{1}+\chi_{7}$ of degree 6. These characters are afforded by the following two representations. One is permutatons of projective coordinates $[x_{1}:x_{2}:x_{3}:x_{4}:x_{5}:x_{6}]$ denoted by $\rho$, and another is the composition of $\rho$ and an outer automorphism $\sigma$ of $\mathrm{A}_{6}$ (see Appendix). We have $V_{3}(\rho)=V_{3}(\rho\circ\sigma)$ (see Proposition 2.6) and this vector space is given by
@@
This linear system contains the Fermat cubic, and a general menber is smooth. Since $\mathrm{A}_{6}$ is a non-commutative simple group, any action is symplectic (Although $\mathrm{S}_{6}$ acts on these cubics, the action of $(12)\in\mathrm{S}_{6}$ is not symplectic since it is diagonalized into $\mathrm{diag}(1,1,1,1,1,-1)$, see Table 2 ). We have $\dim\mathrm{C}_{\rho}(3.\mathrm{A}_{6})=1^{2}+1^{2}=2$ (see Remark 2.8) and
@@
This linear system contains a smooth cubic fourfolds $f(x)=0$, and a general menber is smooth. Moreover we have $H=\overline{H}$ since
@@ 3.2. The dihedral group $\mathrm{D}_{8}$
$\begin{array}[]{c|rrrrr}\hline\cr\rm class&1&\rm 2A&\rm 2B&\rm 2C&\rm 4A\cr&\rm 1&a^{2}&\{b,ba^{2}\}&\{ba,ba^{3}\}&\{a,a^{3}\}\cr\hline\cr\chi_{1}&1&1&1&1&1\cr\chi_{2}&1&1&-1&-1&1\cr\chi_{3}&1&1&1&-1&-1\cr\chi_{4}&1&1&-1&1&-1\cr\chi_{5}&2&-2&0&0&0\cr\hline\cr\end{array}.$
@@
It interchanges classes 2B and 2C, and hence $\chi_{3}$ and $\chi_{4}$. Therefore invariants for $\theta_{3}$ and $\theta_{4}$ are same. Let us construct representatoins $\rho_{i}$ affording $\theta_{i}\ (i=1,2,3)$.

### v2

# arXiv:2409.08448v2[math.AG] 02 Mar 2026

@@ 1. Introduction
More recently, Yang, Yu and Zhu studied full automorphism groups for cubic fivefolds and fourfolds in [YYZ23]. According to them, there are 15 maximal groups, and a finite group $G$ can acts faithfully on a smooth cubic fourfold if and only if $G$ is a subgroup of these 15 groups. However, explicit equations of cubic fourfolds with automorphisms are still unknown in many cases.
In this paper, we study symplectic actions of 34 groups of Laza and Zheng, and detertmine invariant smooth cubic fourfolds. Since any automorphism of a smooth cubic fourfold is given by a projective transformation, our problem is reduced to classification of faithfull projective representations of 34 groups which satisfy symplectic conditions. We classify such projective representations by considering character tables of Schur covers for 34 groups. A Schur cover $G^{*}$ of a finite group $G$ is defined as a central extension $G^{*}=M.G$ such that $M\cong\mathrm{H}^{2}(G,\mathbb{C}^{\times})$ and $M\subset[G^{*},G^{*}]$. It is known that complex projective representations of a finite group $G$ lift to linear representations of a Schur cover $G^{*}$. Therefore, for a fixed $G$, we can obtain all families of cubic fourfolds such that $G\subset\mathrm{Aut}^{s}(X)$ in the following manner.
(1) Determine linear representations $\rho:G^{*}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ which induces a faithful projective action of $G$ and satisfies symplectic conditions.
(2) Check that the vector space of $\rho(G)$-invariant cubic forms $V_{3}(\rho)\subset\mathbb{C}[x_{1},\cdots,x_{6}]$ contains $F(x)\in V_{3}(\rho)$ which gives a smooth cubic fourfold $X_{F}$.
(3) Check that $\mathrm{Aut}^{s}(X_{F})=G$ for a general member $F\in V_{3}(\rho)$.
For (1), we can determine possible representations from character tables relying on absolute values of characters (see section 2.2). We can check (3) by counting dimensions of moduli spaces (see section 2.4). For each representation $\rho$ obtained in this manner, $V_{3}(\rho)$ defines an irreducible family of cubic fourfolds such that $G\subset\mathrm{Aut}^{s}(X)$. However, we may have $V_{3}(\rho_{1})=V_{3}(\rho_{2})$ even if $\rho_{1}$ is not projectively equivalent to $\rho_{2}$. This is caused by the action of $\mathrm{Out}(G)$. For an outer automorphism $\sigma$ of $G$, we may have $\rho\neq\rho\circ\sigma$ although we have $V_{3}(\rho)=V_{3}(\rho\circ\sigma)$. Excluding duplicated families, we have
@@
$3,\ \mathrm{A}_{3,3},\ \mathrm{D}_{12},\ \mathrm{A}_{4},\ \mathrm{A}_{4,3},\ \mathrm{A}_{5},\ 3^{2}.4,\ \mathrm{A}_{6},\ \mathrm{S}_{5},\ \mathrm{A}_{7},\ \mathrm{M}_{10}.$
For other cases, the moduli spaces are irreducible (see Table 2 in Appendix B ).
For maximal cases, defining equations of corresponding cubic fourfolds are known, and they are defined over $\mathbb{Q}$ in most cases. In the last section, we show
@@
Acknowledgements. The author is grateful to referees for many suggestions to improve the manuscript. It was raised by referees that the number of irreducible components is (perhaps more easily) computed by computing the number of primitive embeddings of the lattice $S_{G}=(H^{4}(X,\mathbb{Z})_{prim})_{G}$ into $H^{4}(X,\mathbb{Z})_{prim}$. The number of embeddings is included in [MM24], Table 4. The groups of Laza and Zheng are included in this list, and each time they appear with a $\times$ in the cubic column corresponds to the number of primitive embeddings of $S_{G}$.
@@
$1\longrightarrow A\overset{\iota}{\longrightarrow}G\overset{\pi}{\longrightarrow}B\longrightarrow 1$
@@
  • $p^{1+2n}_{\pm}$ denote extraspecial groups of order $p^{1+2n}$ (see Remark 6.1).
@@ 2.1. Schur covering groups
and the symmetric group $\mathrm{S}_{7}$ acts on it as permutations of coordinates (However, the action of $(12)\in\mathrm{S}_{7}$ is not symplectic). Another $\mathrm{A}_{7}$-invariant cubic fourfold was identified by Yang, Yu and Zhu in [YYZ23] considering the “triple cover” of $\mathrm{A}_{7}$. Their model is defined over $\mathbb{Q}(\sqrt{15})$, and we give an equation over $\mathbb{Q}$ here. Let $G$ be a subgroup of $\mathrm{GL}_{6}(\mathbb{C})$ generated by two matrices
@@
(2.4) $1\longrightarrow M\overset{\iota}{\longrightarrow}G^{*}\longrightarrow G\longrightarrow 1$
such that $M\cong\mathrm{H}^{2}(G,\mathbb{C}^{\times})$ and $\iota(M)\subset[G^{*},G^{*}]$. In general, $G^{*}$ is not uniquely determined. The Klein’s four group $\mathrm{C}_{2}\times\mathrm{C}_{2}$ has two Schur covering groups $\mathrm{D}_{8}$ and $\mathrm{Q}_{8}$. However, it is known that all projective representations of $G$ can be lifted to linear representations of one Schur covering group $G^{*}$. The abelian group $M=M(G)$ in equation (2.4) is called the Schur multiplier of $G$. For example, we have
@@
(see [Ka85]). If $M(G)=1$, we have $G^{*}=G$ and any projective representation of $G$ is lifted to a linear representation of $G$ itself. For non-maximal cases of Laza and Zheng’s 34 groups, Schur covers are given in Appendix B.
@@
It is known that $G$ has only one Schur covering group if the orders of $G/[G,G]$ and $M(G)$ are coprime ([Ka85], Theorem 4.4).
@@ 2.2. Symplectic automorphisms of cubic 4-folds
$1,2,3,4,5,6,7,8,9,11,12,15.$
If a cubic fourfold $X$ has a symplectic automorphism of order $n=9,11,12,15$, then we have
@@
[…]
@@ 2.3. Invariants
If $\chi$ is trivial, we say that $F(x)$ is a $\rho(G)$-invariant. We denote the vector space of $\rho(G)$-semi-invariants of degree $d$ with character $\chi$, by $V_{d}(\rho,\chi)$. If $\chi$ is trivial, it is denoted by $V_{d}(\rho)$.
@@
(3) We have Molien’s formula (see [DK15])
$\frac{1}{|G|}\sum_{g\in G}\frac{1}{\det(\mathbf{1}_{n}-t\rho(g))}=\sum_{d=0}^{\infty}(\dim V_{d}(\rho))t^{d},$
and we can compute $\dim V_{d}(\rho)$ using a function “MolienSeries” in GAP.
@@
[…]
@@ 2.4. Moduli spaces
$\mathrm{C}_{\rho}(G)$ $=\{h\in\mathrm{GL}_{n}(\mathbb{C})\ |\ hgh^{-1}=g\ \ \text{for}\ \ {}^{\forall}g\in\rho(G)\},$
$\mathrm{N}_{\rho}(G)$ $=\{h\in\mathrm{GL}_{n}(\mathbb{C})\ |\ h\rho(G)h^{-1}=\rho(G)\}.$
@@
$\mathrm{N}_{\rho,\chi}(G)$ $=\{h\in\mathrm{N}_{\rho}(G)\ |\ \chi_{\rho}(hgh^{-1})=\chi_{\rho}(g)\ \ \text{for}\ \ {}^{\forall}g\in\rho(G)\}$
@@
Now let $G$ be one of Laza and Zheng’s 34 groups, $G^{*}$ be a Schur cover and $\rho:G^{*}\rightarrow\mathrm{GL}_{6}(\mathbb{C})$ be a lifting of a faithful projective action of $G$ on $\mathbb{P}^{5}$. If $\rho(G^{*})$ acts simplectically on a smooth cubic fourfold $F(x)=0$ with $F\in V_{3}(\rho,\,\chi)$, we have
@@
Note that we have $\dim\mathrm{N}_{\rho,\chi}(G^{*})\geq 1$ since $\mathbb{C}^{\times}\subset\mathrm{N}_{\rho,\chi}(G^{*})$. Therefore we have the following criterion to exclude unnecessary cases.
@@ 2.5. $\mathrm{A}_{6}$-invariant cubic fourfolds
[…]
@@
Classes 1, 3A and 3B form the center $\mathrm{C}_{3}$, and $\chi_{i}\ (i=1,6,7)$ are irreducible characters with $\mathrm{C}_{3}\subset\mathrm{Ker}\,\chi_{i}$, that is, characters of $\mathrm{A}_{6}$ itself rather than the triple cover $3.\mathrm{A}_{6}$. In this case, we have two (non-trivial) characters $\chi_{1}+\chi_{6}$ and $\chi_{1}+\chi_{7}$ of degree 6. These characters are afforded by the following two representations. One is permutatons of projective coordinates $x_{1}:x_{2}:x_{3}:x_{4}:x_{5}:x_{6}]$ denoted by $\rho$, and another is the composition of $\rho$ and an outer automorphism $\sigma$ of $\mathrm{A}_{6}$ (see Appendix A). We have $V_{3}(\rho)=V_{3}(\rho\circ\sigma)$ (see Proposition 2.6) and this vector space is given by
@@
This linear system contains the Fermat cubic, and a general member is smooth. Since $\mathrm{A}_{6}$ is a non-commutative simple group, any action is symplectic (Although $\mathrm{S}_{6}$ acts on these cubics, the action of $(12)\in\mathrm{S}_{6}$ is not symplectic since it is diagonalized into $\mathrm{diag}(1,1,1,1,1,-1)$, see Table 1 ). We have $\dim\mathrm{C}_{\rho}(3.\mathrm{A}_{6})=1^{2}+1^{2}=2$ (see Remark 2.8) and
@@
This linear system contains a smooth cubic fourfolds $f(x)=0$, and a general member is smooth. Moreover we have $H=\overline{H}$ since
@@
2.6. Proof of Theorem 1.1
As is the case of $\mathrm{A}_{6}$, we determine all admisible projective representations (excluding $\rho$ for which $V_{3}(\rho)$ is same with invariants already obtained) for 34 groups $G$ in the following sections. The result is listed in Table 2 in Appendix B. To complete the proof of Theorem 1.1, we need to show that two families for groups
$3,\ \mathrm{A}_{3,3},\ \mathrm{D}_{12},\ \mathrm{A}_{4},\ \mathrm{A}_{4,3},\ \mathrm{A}_{5},\ 3^{2}.4,\ \mathrm{A}_{6},\ \mathrm{S}_{5},\ \mathrm{A}_{7},\ \mathrm{M}_{10}$
[…]
where horizontal arrows are group actions. Therefore if two family $V_{3}(\rho_{1},\chi_{1})$ and $V_{3}(\rho_{2},\chi_{2})$ for $G$ listed in the above are same, then there exist $\phi\in\mathrm{Aut}(G)$ such that $\rho_{1}$ is projectively equivalent to $\rho_{2}\circ\phi$. However, we see that this is impossible by considering actions of Schur multiplier or comparing dimensions of irreducible components of two representations. For example, we have two families (2.6) and (2.9) for $G=\mathrm{A}_{6}$. We can distinguish them from each other since the Schur multiplier $\mathrm{C}_{3}$ acts trvially for (2.6) and non-trivially for (2.9). We see also that irreducible decomposition of the action of $\mathrm{A}_{6}$ consists of a 1-dimensional representation and a 5-dimensional representatoin for (2.6) but an irreducible 6-dimensional representation for (2.9).
@@ 3.2. The dihedral group $\mathrm{D}_{8}$
$\begin{array}[]{c|rrrrr}\hline\cr\rm class&1&\rm 2A&\rm 2B\quad&\rm 2C\quad&\rm 4A\quad\cr&\rm 1&a^{2}&\{b,ba^{2}\}&\{ba,ba^{3}\}&\{a,a^{3}\}\cr\hline\cr\chi_{1}&1&1&1\quad&1\quad&1\quad\cr\chi_{2}&1&1&-1\quad&-1\quad&1\quad\cr\chi_{3}&1&1&1\quad&-1\quad&-1\quad\cr\chi_{4}&1&1&-1\quad&1\quad&-1\quad\cr\chi_{5}&2&-2&0\quad&0\quad&0\quad\cr\hline\cr\end{array}.$
@@
It interchanges classes 2B and 2C, and hence $\chi_{3}$ and $\chi_{4}$. Therefore invariants for $\theta_{3}$ and $\theta_{4}$ are same. Let us construct representations $\rho_{i}$ affording $\theta_{i}\ (i=1,2,3)$.